Suppose the market demand and supply function for cookies per day in Puri city is given as Qd = 20-2P and Qs = - 8+2P, where Qd, Qs and P refers to quantity demanded, quantity supplied and price respectively. Draw the demand and supply graph and show the market equilibrium price and quantity. Analyse the market situation if the actual price of cookies is Rs. 10 per unit by drawing a well labeled graph.


1
Expert's answer
2021-03-01T07:27:46-0500

(a)

Qs=Qd

"-8+2p=20-2p"

"2p+2p=20+8"

"4p=28"

"P=7"

"Q=6"



(b)

According to the law of demand and supply an increase in price will lead to a decrease in demand and increase in supply.

Therefore the demand for cookies will decrease as their supply increases.